OpenStudy (anonymous):

Please help what is 3/8 + 1/6?

OpenStudy (anonymous):

\[\dfrac{3}{8} + \dfrac{1}{6} = \dfrac{9}{24} + \dfrac{4}{24}\]

OpenStudy (anonymous):

3/8 + 1/6. U HAVE TO FIND THE LEAST COMMON DENOMINATOR